OpenStudy (anonymous):

Use the quadratic formula to find any x-intercepts on the graph of the equation. y=x^2-2x-4

OpenStudy (anonymous):

first, identify the a, b, c in your equation. to find the roots (x-intercepts) of \(\large y=ax^2+bx+c \), the quadratic formula is: \(\large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a} \)

OpenStudy (anonymous):

I have that I just do not know how to solve it.

OpenStudy (anonymous):

A = 1 B=-2 C=-4

OpenStudy (anonymous):

good... now just stick those values into the formula and simplify......

OpenStudy (anonymous):

\(\huge x=\frac{-(-2) \pm \sqrt{(-2)^2-4(1)(-4)}}{2(1)} \)

OpenStudy (anonymous):

Right I have all that and i have it down to -2√24/2

OpenStudy (anonymous):

Reduced it would be -2 \[\sqrt{12}\]

OpenStudy (precal):